OpenStudy (anonymous):

simplify 56 7/8 + 3/5 + 1/8 + 4 2/5

OpenStudy (anonymous):

have you attempted to answer?

OpenStudy (anonymous):

no i have no idea

OpenStudy (anonymous):

what is 7/8+1/8

OpenStudy (anonymous):

8/8? which is 1

OpenStudy (anonymous):

so 56 7/8 +1/8 =

OpenStudy (anonymous):

57?

OpenStudy (anonymous):

@rayleigh you're right, so in the same manner, what is 4 2/5 +3/5=

OpenStudy (anonymous):

Simplify the first mixed fraction to improper fraction and then add all fractions with demo. 8 on one side and all fractions with deno. 5 , you will get the numerator completely divisible by denominator and answer as 62

OpenStudy (anonymous):

5

OpenStudy (anonymous):

@rayleigh right again, now you have simplified to 57+5 =

OpenStudy (anonymous):

I hope you got the answer

OpenStudy (anonymous):

which is 62

OpenStudy (anonymous):

@rayleigh right again, remember the simplest solution to a problem is usually the correct one!

OpenStudy (anonymous):

alright thank you